OpenStudy (help!!!!):

What are the equation for conservation of energy equations for gravitational, elastic, chemical, kinetic energy?

OpenStudy (sachin00007):

Intitial energy +work done = final energy Kinetic energy = 1/2 mv^2 G potential energy= mgh

OpenStudy (sachin00007):

Elastic potential energy = 1/2 ke^2 Where k is the spring constant And e is the extension